Linear Equations in one variable, Solve for x:
50-x-(3x+2)=0

Answer by matthew_sessoms(39)

Distribute the negative (-) for the parentheses.


50 - x - 3x - 2 = 0


Combine like terms...
48 - 4x = 0
48 = 4x
12 = x


Check
50 - 12 -(3(12) + 2) = 0
38 - (38) = 0
0 = 0
TRUE


-Matthew